Sanjana loves to read books(B) and likes to watch movies(M) as well. She derives double the satisfaction from reading a book than watching a movie. Her utility function is as follows

U(B, M) = 2B + M

Where B and M are the number of books and movies respectively.

a. Explain the shape of the indifference curve using diagrams. Word limit: 200]

b. The income available with her is INR60. And price of movie ticket is INR 10 and that of a book is INR 15. Will she spend her entire budget and why? What will be her consumption, explain why?


1
Expert's answer
2021-12-05T18:48:38-0500

a. The indifference curve in this case can be shown as a downward-sloping line.

b. She will she spend her entire budget at point, for which MUb/Pb = MUm/Pm.

MUb = U'(b) = 2,

MUm = U'(M) = 1,

2/1 = 15/10 is not possible, so at current prices her utility is not maximized.